# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 7. New Examples of $\mathsf{RCD}$-spaces of cohomogeneity one
We observe that $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is the homogeneous space $\mathrm{SO}(3)\times\mathrm{SO}(3)/(\mathrm{SO}(2)\times\mathrm{SO}(2))\cong(\mathrm{SO}(3)/\mathrm{SO}(2))\times(\mathrm{SO}(3)/\mathrm{SO}(2))$. Thus by Theorem 7.1, there exists a $\mathrm{SO}(3)\times\mathrm{SO}(3)$-invariant Riemannian metric $g$ on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ such that for any $K\geq 0$ the $(K,5)$-cone $\mathrm{Con}^{5}_{K}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ is a non-collapsed $\mathsf{RCD}(K4,5)$-space with a cohomogeneity one action by measure preserving isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$. But for $K>0$ the $(K,5)$-cone is homeomorphic to the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$. This proves the first claim.
Assume that the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits an Alexandrov metric with an action of cohomogeneity one by isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ fixing the vertices of the suspension and with principal isotropy $\mathrm{SO}(2)\times\mathrm{SO}(2)$. Consider $X$ the space of directions at one of the vertexes of the suspension. This space is homeomorphic to $\mathbf{S}^{2}\times\mathbf{S}^{2}$, $\mathrm{SO}(3)\times\mathrm{SO}(3)$ acts transitively on $X$, and it is a $4$-dimensional Alexandrov space of curvature at least one. Thus $X$ is a homogeneous space, and by [5, Theorem 7 I] $X$ is isometric to a homogeneous Finsler manifold. Since $X$ is an Alexandrov space of curvature at least $1$, then it is also an $\mathsf{RCD}(3,4)$-space, then $X$ is isometric to a homogenous Riemannian manifold $(M,g)$. Moreover we have that $\sec(g)\geq 1$. Since we have non-trivial Killing vector fields by the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$, by [50, Theorem 1] $M$ is homeomorphic to $\mathbf{S}^{2}$ or $\mathbb{C}P^{2}$, which is a contradiction. Thus $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is not an $\mathrm{SO}(3)\times\mathrm{SO}(3)$-cohomogeneity one Alexandrov space. ∎
@@
When considering $G/H$ a homogeneous smooth manifold with finite fundamental group, using Theorem B for the group diagram $(G,H,G,G)$ we also get an $\mathsf{RCD}(K,N)$-space. From the proof if follows that this case $K=0$ and $N=\dim(G/H)+\dim(G)+1$, and the orbit space is isometric to $[-1-2/A,1+2/A]$ for an appropriate choice of $A$ and measure $f^{\dim(G/H)}\mathcal{L}$, where $f$ is a non-smooth function. Nonetheless, this space is equivariantly homeomorphic to the suspension of $G/H$ and thus to the space given by Theorem 7.1. That is, for a group diagram $(G,H,G,G)$ with $G/H$ having finite fundamental group, by Theorem B and Theorem 7.1 we obtain two equivariantly homeomorphic spaces, but each one with a different $\mathsf{RCD}$-structure.

### v3

@@ 7. New Examples of $\mathsf{RCD}$-spaces of cohomogeneity one
We observe that $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is the homogeneous space $\mathrm{SO}(3)\times\mathrm{SO}(3)/(\mathrm{SO}(2)\times\mathrm{SO}(2))\cong(\mathrm{SO}(3)/\mathrm{SO}(2))\times(\mathrm{SO}(3)/\mathrm{SO}(2))$. Thus by ?THM? LABEL:T:_Suspensions_are_RCD_spaces_of_coho_1, there exists a $\mathrm{SO}(3)\times\mathrm{SO}(3)$-invariant Riemannian metric $g$ on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ such that for any $K\geq 0$ the $(K,5)$-cone $\mathrm{Con}^{5}_{K}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ is a non-collapsed $\mathsf{RCD}(K,5)$-space with a cohomogeneity one action by measure preserving isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$. But for $K>0$ the $(K,5)$-cone is homeomorphic to the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$. This proves the first claim.
Assume that the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits an Alexandrov metric with an action of cohomogeneity one by isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ fixing the vertices of the suspension and with principal isotropy $\mathrm{SO}(2)\times\mathrm{SO}(2)$. Consider $X$ the space of directions at one of the vertexes of the suspension. This space is homeomorphic to $\mathbf{S}^{2}\times\mathbf{S}^{2}$, $\mathrm{SO}(3)\times\mathrm{SO}(3)$ acts transitively on $X$, and it is a $4$-dimensional Alexandrov space of curvature at least one. Thus $X$ is a homogeneous space, and by [8, Theorem 7 I] $X$ is isometric to a homogeneous Finsler manifold. Since $X$ is an Alexandrov space of curvature at least $1$, then it is also an $\mathsf{RCD}(3,4)$-space, then $X$ is isometric to a homogeneous Riemannian manifold $(M,g)$. Moreover we have that $\sec(g)\geq 1$. Since we have non-trivial Killing vector fields by the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$, by [49, Theorem 1] $M$ is homeomorphic to $\mathbf{S}^{4}$ or $\mathbb{C}P^{2}$, which is a contradiction. Thus $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is not an $\mathrm{SO}(3)\times\mathrm{SO}(3)$-cohomogeneity one Alexandrov space. ∎
@@
When considering $G/H$ a homogeneous smooth manifold with finite fundamental group, using ?THM? LABEL:thm.rcd_space_from_group_diagram for the group diagram $(G,H,G,G)$ we also get an $\mathsf{RCD}(K,N)$-space. From the proof if follows that this case $K=0$ and $N=\dim(G/H)+\dim(G)+1$, and the orbit space is isometric to [12/A,1+2/A][-1-2/A,1+2/A] for an appropriate choice of $A$ and measure $f^{\dim(G/H)}\mathcal{L}$, where $f$ is a non-smooth function. Nonetheless, this space is equivariantly homeomorphic to the suspension of $G/H$ and thus to the space given by ?THM? LABEL:T:_Suspensions_are_RCD_spaces_of_coho_1. That is, for a group diagram $(G,H,G,G)$ with $G/H$ having finite fundamental group, by ?THM? LABEL:thm.rcd_space_from_group_diagram and ?THM? LABEL:T:_Suspensions_are_RCD_spaces_of_coho_1 we obtain two equivariantly homeomorphic spaces, but each one with a different $\mathsf{RCD}$-structure.